OpenStudy (anonymous):

y=6x^2+12x+10 1. Change from polynomial form to vertex form 2.does it face up or down 3.vertex 4. y interecept 5. line of symmetery 6. narrow, wide, or parent? 7. is it a min or max help please? i have to give a presentation tomorrow :(

OpenStudy (anonymous):

Hey for this question Do you know how to complete the square?

OpenStudy (anonymous):

that where im stuck i have the way it faces, min and line of symmetry but I cant seem to complete the square properly

OpenStudy (anonymous):

Give me a second this i get this weird answer let me try again

OpenStudy (anonymous):

So do you want me to tell you how to complete it first?

OpenStudy (anonymous):

sure that would be great :)

OpenStudy (anonymous):

y=6x^2+12x+10 0=6x^2+12x+10 Divde the 6 to all of the numbers x^2+2x +10/6= 0 Simplify x^2 +2x+ 5/3 =0 Move the 5/3 to other side by subtracting it. x^2 +2x + (Box)= 5/3 +(box) x^2 +2x + (1) = 5/3 +(1) (x+1)^2 = 8/3 5/3+1 =5/3 +3/3 Square root both sides x+1= square root of 8/3 subtract 1 which is x= 1+ or - square root of 8/3

OpenStudy (anonymous):

Sorry its really long haha

OpenStudy (anonymous):

that confused me one sec let me think about it haha

OpenStudy (anonymous):

Haha Alright but i have to go soon :p

OpenStudy (anonymous):

Hope i did it right.

OpenStudy (anonymous):

i did it like this 6x^2+12x+10 2(3x^2+6x )+5 2(3x^2+6x+1)+5-2 and thats where i got losst

OpenStudy (anonymous):

Oh i see on completing square you cannot factor it like that.

OpenStudy (anonymous):

You have to completely divide it :( http://www.purplemath.com/modules/sqrquad.htm Check it out the first question with 4x

OpenStudy (anonymous):

oh dang thats how my teacher t aught me :(

OpenStudy (anonymous):

So you got the rest down?

OpenStudy (anonymous):

yeah once i get it i can finsh soslving because ill know h and k thank you :)

OpenStudy (anonymous):

haha no problem i hope you do well and just in case please check it over :( i dont want you to get bad grade because of my mistake

OpenStudy (anonymous):

oh dont worry about it haha youve helped me i would have been completely lost lol

OpenStudy (anonymous):

Haha Good luck :p good night :3

OpenStudy (anonymous):

thanks :) good night